If sin A = cos A
find A
1
Expert's answer
2012-06-19T10:47:47-0400
First of all, let's substitute

cos x = sin (x + pi/2).

This gives

sin x = sin (x + pi/2),

wheach leads to x = pi/4 ± pi·k, where k is some integer.